# A martingale approach to noncommutative stochastic calculus

### v2

@@ 1 Introduction
whenever $X$ is a (continuous vector-valued) semimartingale and $F$ is a $C^{2}$ map; above, $DF$ and $D^{2}F$ are, respectively, the first and second Fréchet derivatives of $F$, and $\mathrm{d}X_{s},\mathrm{d}X_{s}]$ denotes integration against the quadratic variation of $X$. This crucial perspective from classical stochastic analysis has not yet made it into the noncommutative probability literature. The central goal of the present paper is to incorporate this insight by developing a general theory of noncommutative stochastic calculus that follows as closely as possible the classical martingale-theoretic development of the subject. As we summarize in more detail in Section 1.2, the fruits of these labors are the first general theory of noncommutative quadratic (co)variation, continuous-time noncommutative Bukholder–Davis–Gundy inequalities (Theorem 1.11), and a noncommutative Itô’s formula (Theorem 1.13) in which the correction term is a quadratic variation integral of the second derivative as in the classical case. Moreover, we show that other known instances of noncommutative Itô’s formula arise as special cases of ours via derivative and quadratic variation calculations, thereby demonstrating a “universality” of our formula. A key point is that we consider general maps $F$ defined on subsets of the operator algebra rather than highly specific classes of functions like those induced via functional calculus by a scalar function of a real variable.
@@
In Section 4, we develop a general theory of stochastic integration with respect to the $L^{2}$-decomposable processes introduced in Section 3.2. We do so by adapting the classical method, explained in Section 1.1, of proving an Itô isometry using the Doléans measure of a square-integrable martingale. In Section 5, the most technically challenging part, we develop our theory of noncommutative quadratic covariation. In Section 5.4, we obtain our noncommutative Bukholder–Davis–Gundy inequalities. We also compute examples of quadratic covariations in Section 5.5 that shed new light on other calculations in the literature.
@@ 1.1 Philosophy of the approach
Suppose $\infty<a<b\leq\infty$, and write $I\coloneqq[a,b]\cap\mathbb{R}$. If $b<\infty$, then a partition of $I$ is a finite subset $\Pi=\{a=t_{0}<\cdots<t_{n}=b\}\subseteq I$. A partition of $a,\infty)$ is a collection $\Pi=\{t_{n}:n\in\mathbb{N}_{0}\}$ such that $t_{0}=a$, $t_{n}<t_{n+1}$ for all $n\in\mathbb{N}_{0}$, and $t_{n}\to\infty$ as $n\to\infty$. In general, $\mathcal{P}_{I}$ is the set of partitions of $I$. Now, fix $\Pi\in\mathcal{P}_{I}$. If $t\in\Pi$, then $t_{-}\in\Pi$ is the member of $\Pi$ to the left of $t$; precisely, $a_{-}\coloneqq a$ and $t_{-}\coloneqq\max\{s\in\Pi:s<t\}$ for $t\in\Pi\setminus\{a\}$. Also, $\Delta t\coloneqq t-t_{-}$, $|\Pi|\coloneqq\sup\{\Delta s:s\in\Pi\}$ is the mesh of $\Pi$, and $\Delta_{t}F\coloneqq F(t)-F(t_{-})$ for a function $F$ from $I$ to a vector space. Limits as $|\Pi|\to 0$ will be denoted by $\lim_{\Pi\in\mathcal{P}_{I}}$; see Fact 4.2 below.
@@
This allows for the extension of the definition of $\int_{0}^{\boldsymbol{\cdot}}H_{s}\,\mathrm{d}M_{s}\coloneqq I_{M}(H)$ to the set of predictable, i.e., $\mathcal{P}$-measurable, processes $H\colon\mathbb{R}_{+}\times\Omega\to\mathbb{R}$ such that $\int_{[0,t]\times\Omega}|H|^{2}\,\mathrm{d}\mu_{M}<\infty$ for all $t\geq 0$. (See [14, §2.5] for details.) Finally, if $H$ is bounded, continuous, and adapted, then the Itô isometry and the approximation $H_{r}^{\Pi}\coloneqq\sum_{s\in\Pi}1_{(s_{-},s]}(r)\,H_{s_{-}}$ can be used to show that
@@
    For each pair $(X,Y)$ of “noncommutative semimartingales” and sufficiently many “adapted, bilinear map–valued processes $\Lambda$,” construct a quadratic covariation integral $\int_{0}^{\boldsymbol{\cdot}}\Lambda(t)[\mathrm{d}X(t),\mathrm{d}Y(t)]$, preferably in such a way that (1.2) has a noncommutative analog.
@@ 1.2 Summary of main results
To begin, let $p\in[1,\infty]$. A process $X\colon\mathbb{R}_{+}\to L^{p}(\mathcal{A},\mathbb{E})=L^{p}(\mathbb{E})$ is adapted if $X(t)\in L^{p}(\mathcal{A}_{t},\mathbb{E})$ for all $t\geq 0$. An $\boldsymbol{L^{p}}$-FV process is an adapted process $X\colon\mathbb{R}_{+}\to L^{p}(\mathbb{E})$ with locally bounded variation with respect to the noncommutative $L^{p}$ norm. By analogy with the classical notion of a semimartingale, we consider processes that can be decomposed as the sum of a martingale and an FV process: $X$ is a (continuous) $\boldsymbol{L^{p}}$-decomposable process if $X=X(0)+M+A$ for some $L^{p}$-continuous martingale $M\colon\mathbb{R}_{+}\to L^{p}(\mathbb{E})$ (Definition 3.11(ii)) and some $L^{p}$-continuous $L^{p}$-FV process $A\colon\mathbb{R}_{+}\to L^{p}(\mathbb{E})$ with $M(0)=A(0)=0$. In Section 3.2, we give many examples of $L^{p}$-decomposable processes and show that if $p\geq 2$, then the decomposition $X=X(0)+M+A$ is unique (Corollary 3.22). As in the classical case, we call $M$ the martingale part of $X$ and $A$ the FV part of $X$.
@@
We are now ready to state our first main result: a noncommutative analog of Theorem 1.2.
@@
whenever $A,B\colon\mathbb{R}_{+}\to\mathcal{A}$ are adapted and continuous. These are special cases of integrals of trace biprocesses (Definition 3.7), which we introduce at the end of Section 3.1. The term “trace biprocess” is inspired by previous work on noncommutative stochastic calculus—specifically, that of Biane–Speicher [8]—in which integrals like $\int_{0}^{\boldsymbol{\cdot}}A(t)\,\mathrm{d}X(t)\,B(t)$ are treated by defining integrals of tensor-valued process called “biprocesses.” Notably, however, integrals like $\int_{0}^{\boldsymbol{\cdot}}\mathbb{E}[A(t)\,\mathrm{d}X(t)]\,B(t)$ are not considered in previous work.
@@
Here is the second main result: a construction of noncommutative quadratic covariation.
@@
Furthermore, $\lVert M(t)\rVert_{2}^{2}=\mathbb{E}\big{[}\int_{0}^{t}\mathrm{d}M^{*}(s)\,\mathrm{d}M(s)\big{]}=\mathbb{E}\big{[}\int_{0}^{t}\mathrm{d}M(s)\,\mathrm{d}M^{*}(s)\big{]}$.
@@
If $d,\ell\in\mathbb{N}$ and $f\colon\mathbb{R}^{d}\to\mathbb{C}^{\ell}$ is $k$-times continuously differentiable and $f_{*}(\mathbf{a})\coloneqq f\circ\mathbf{a}\in\mathcal{A}^{\ell}$ for all $\mathbf{a}\in\mathcal{A}_{\operatorname{sa}}^{d}$, then $f_{*}\colon\mathcal{A}_{\operatorname{sa}}^{d}\to\mathcal{A}^{\ell}$ is an adapted $C^{k}$ map (Proposition 6.9 with $n=m=1$).
@@ 2.1 Noncommutative probability
$(L^{p}(\tau_{n}),\lVert\cdot\rVert_{L^{p}(\tau_{n})})=\Big{(}L^{p}(\Omega,\mathscr{F},P;\mathrm{M}_{n}(\mathbb{C})),\big{\|}\lVert\cdot\rVert_{L^{p}(\operatorname{tr}_{n})}\big{\|}_{L^{p}(\mathbb{E}_{P})}\Big{)}\qquad(p\in[1,\infty]).$
@@ 2.2 Multilinear maps on $L^{p}$ spaces
We shall omit the subscripts when $k=1$. Finally, define $\mathbb{B}_{0}\coloneqq\mathcal{A}_{0+1}=\mathcal{A}_{1}$ for the $k=0$ case.
@@ 2.3 Trace $\ast$-polynomials
For example, $P(x_{1},x_{2},x_{3})=\operatorname{tr}(x_{1}x_{2}x_{1})x_{3}x_{2}^{2}x_{3}-\operatorname{tr}(x_{1})\operatorname{tr}(x_{2})x_{3}$ belongs to $\mathrm{TrP}(x_{1},x_{2},x_{3})$, and we also have that $P(x_{1},x_{2},x_{3})=\operatorname{tr}(x_{1}^{2}x_{2})x_{3}x_{2}^{2}x_{3}-\operatorname{tr}(x_{1})\operatorname{tr}(x_{2})x_{3}$ because $\operatorname{tr}$ is “tracial.”
@@
  • $y_{j}^{*}$ appears precisely once in $P_{0}\cdots P_{\ell}$, and and $y_{j}$ does not appear in $P_{0}\cdots P_{\ell}$.
@@ 3.1 Filtrations and adaptedness
For the remainder of this section, write $(\mathcal{A},(\mathcal{A}_{t})_{t\geq 0},\mathbb{E})\coloneqq(\mathcal{A}_{1},(\mathcal{A}_{1,t})_{t\geq 0},\mathbb{E}_{1})$, and fix $k\in\mathbb{N}$. All tensor products below are over $\mathbb{C}$.
@@
    $_{k}(a_{1}\otimes\cdots\otimes a_{k+1})[b_{1},\ldots,b_{k}]=a_{1}b_{1}\cdots a_{k}b_{k}a_{k+1}\qquad(a_{1},\ldots,a_{k+1},b_{1},\ldots,b_{k}\in\mathcal{A}).$
@@
    $_{k}^{\mathbb{E}}(a_{1}\otimes\cdots\otimes a_{k+1})[b_{1},\ldots,b_{k}]=\mathbb{E}[a_{1}b_{1}\cdots a_{k}b_{k}]a_{k+1}\qquad(a_{1},\ldots,a_{k+1},b_{1},\ldots,b_{k}\in\mathcal{A}).$
@@ 4.1 Stieltjes integrals
Also, if $(\Pi,\xi)\in\mathcal{P}_{I}^{*}$, then
$F^{(\Pi,\xi)}\coloneqq 1_{\{a\}}F(a)+\sum_{s\in\Pi}1_{(s_{-},s]}F(s_{*})\colon I\to\mathcal{V}.$
If $t_{*}=\xi(t)$ is always the left endpoint $t_{-}$ of $t_{-},t]$, then $F^{\Pi}\coloneqq F^{(\Pi,\xi)}$.
The sets $\mathcal{P}_{I}$ and $\mathcal{P}_{I}^{*}$ are frequently directed by refinement. In this paper, we shall instead direct them by mesh $|\cdot|$, i.e., $\Pi\leq\Pi^{\prime}$ and $(\Pi,\xi)\leq(\Pi^{\prime},\xi^{\prime})$ whenever $|\Pi|\geq|\Pi^{\prime}|$. Here is an elementary clarification of what convergence of nets directed by partition mesh means. We leave the proof to the reader.
@@
    If $H$ is continuous and $\mathcal{V}$ is normed, then $H^{(\Pi,\xi)}\to H$ uniformly on compact subsets of $I$ as $|\Pi|\to 0$.
  • Proof.
    As the reader may easily verify, it suffices to treat the $b<\infty$ case. Write $\iota\colon I\to\mathbb{R}$ for the inclusion. Observe that if $(\Pi,\xi)\in\mathcal{P}_{I}^{*}$ and $t\in(a,b]$, then $H^{(\Pi,\xi)}=H\circ\iota^{(\Pi,\xi)}$, $0<t-\iota^{\Pi}(t)\leq|\Pi|$, and $\big{|}t-\iota^{(\Pi,\xi)}(t)\big{|}\leq|\Pi|$. Using these observations, the desired results follow easily from the definitions and, in the case of (ii), the fact that continuous functions $I\to\mathcal{V}$ are uniformly continuous. ∎
@@
$\sum_{t\in\Pi}H(t_{*})[F(t\wedge\cdot)-F(t_{-}\wedge\cdot)]\xrightarrow[(\Pi,\xi)\in\mathcal{P}_{I}^{*}]{|\Pi|\to 0}\int_{a}^{\boldsymbol{\cdot}}H(t)[\mathrm{d}F(t)]$
@@
    In particular, the integral map $\mathbb{S}_{I}\ni S\mapsto\int_{a}^{\cdot}S[\mathrm{d}\mu_{F}^{0}]\in\ell^{\infty}(I;\mathcal{W})$ extends uniquely to a bounded linear map $J\colon\overline{\mathbb{S}_{I}}\to\ell^{\infty}(I;\mathcal{W})$, where $\overline{\mathbb{S}_{I}}$ is the closure of $\mathbb{S}_{I}$ in $\ell^{\infty}(I;B(\mathcal{V};\mathcal{W}))$. Now, if $H\colon I\to B(\mathcal{V};\mathcal{W})$ is continuous and $(\Pi,\xi)\in\mathcal{P}_{I}^{*}$, then $H^{(\Pi,\xi)}\in\mathbb{S}_{I}$, and by Lemma 4.4(ii), $H^{(\Pi,\xi)}\to H$ in $\ell^{\infty}(I;B(\mathcal{V};\mathcal{W}))$ as $|\Pi|\to 0$. Thus, $H\in\overline{\mathbb{S}_{I}}$, and by the continuity of $J$,
    $\sum_{t\in\Pi}H(t_{*})[F(t\wedge\cdot)-F(t_{-}\wedge\cdot)]=\int_{a}^{\cdot}H^{(\Pi,\xi)}\big{[}\mathrm{d}\mu_{F}^{0}\big{]}=J\big{[}H^{(\Pi,\xi)}\big{]}\xrightarrow[(\Pi,\xi)\in\mathcal{P}_{I}^{*}]{|\Pi|\to 0}J[H]$

### v3

@@ 1 Introduction
whenever $X$ is a (continuous vector-valued) semimartingale and $F$ is a $C^{2}$ map; above, $DF$ and $D^{2}F$ are, respectively, the first and second Fréchet derivatives of $F$, and $\mathrm{d}X_{s},\mathrm{d}X_{s}]$ denotes integration against the quadratic variation of $X$. This crucial perspective from classical stochastic analysis has not yet made it into the noncommutative probability literature. The central goal of the present paper is to incorporate this insight by developing a general theory of noncommutative stochastic calculus that follows as closely as possible the classical martingale-theoretic development of the subject. As we summarize in more detail in Section 1.2, the fruits of these labors are the first general theory of noncommutative quadratic (co)variation, continuous-time noncommutative Burkholder–Davis–Gundy inequalities (Theorem 1.11), and a noncommutative Itô’s formula (Theorem 1.13) in which the correction term is a quadratic variation integral of the second derivative as in the classical case. Moreover, we show that other known instances of noncommutative Itô’s formula arise as special cases of ours via derivative and quadratic variation calculations, thereby demonstrating a “universality” of our formula. A key point is that we consider general maps $F$ defined on subsets of the operator algebra rather than highly specific classes of functions like those induced via functional calculus by a scalar function of a real variable.
@@
In Section 4, we develop a general theory of stochastic integration with respect to the $L^{2}$-decomposable processes introduced in Section 3.2. We do so by adapting the classical method, explained in Section 1.1, of proving an Itô isometry using the Doléans measure of a square-integrable martingale. In Section 5, the most technically challenging part, we develop our theory of noncommutative quadratic covariation. In Section 5.4, we obtain our noncommutative Burkholder–Davis–Gundy inequalities. We also compute examples of quadratic covariations in Section 5.5 that shed new light on other calculations in the literature.
@@ 1.1 Philosophy of the approach
Suppose $\infty<a<b\leq\infty$, and write $I\coloneqq[a,b]\cap\mathbb{R}$. If $b<\infty$, then a partition of $I$ is a finite subset $\Pi=\{a=t_{0}<\cdots<t_{n}=b\}\subseteq I$. A partition of $a,\infty)$ is a collection $\Pi=\{t_{n}:n\in\mathbb{N}_{0}\}$ such that $t_{0}=a$, $t_{n}<t_{n+1}$ for all $n\in\mathbb{N}_{0}$, and $t_{n}\to\infty$ as $n\to\infty$. In general, $\mathcal{P}_{I}$ is the set of partitions of $I$. Now, fix $\Pi\in\mathcal{P}_{I}$. If $t\in\Pi$, then $t_{-}\in\Pi$ is the member of $\Pi$ to the left of $t$; precisely, $a_{-}\coloneqq a$, and $t_{-}\coloneqq\max\{s\in\Pi:s<t\}$ for $t\in\Pi\setminus\{a\}$. Also, $\Delta t\coloneqq t-t_{-}$, $|\Pi|\coloneqq\sup\{\Delta s:s\in\Pi\}$ is the mesh of $\Pi$, and $\Delta_{t}F\coloneqq F(t)-F(t_{-})$ for a function $F$ from $I$ to a vector space. Limits as $|\Pi|\to 0$ will be denoted by $\lim_{\Pi\in\mathcal{P}_{I}}$; see Fact 4.2 below.
@@
This enables the extension of the definition of $\int_{0}^{\boldsymbol{\cdot}}H_{s}\,\mathrm{d}M_{s}\coloneqq I_{M}(H)$ to the set of predictable, i.e., $\mathcal{P}$-measurable, processes $H\colon\mathbb{R}_{+}\times\Omega\to\mathbb{R}$ such that $\int_{[0,t]\times\Omega}|H|^{2}\,\mathrm{d}\mu_{M}<\infty$ for all $t\geq 0$. (See [14, §2.5] for details.) Finally, if $H$ is bounded, continuous, and adapted, then the Itô isometry and the approximation $H_{r}^{\Pi}\coloneqq\sum_{s\in\Pi}1_{(s_{-},s]}(r)\,H_{s_{-}}$ can be used to show that
@@
    For each pair $(X,Y)$ of “noncommutative semimartingales” and sufficiently many “adapted, bilinear map–valued processes $\Lambda$,” construct a quadratic covariation integral $\int_{0}^{\boldsymbol{\cdot}}\Lambda(t)[\mathrm{d}X(t),\mathrm{d}Y(t)]$, preferably in a way that (1.2) has a noncommutative analog.
@@ 1.2 Summary of main results
Let $p\in[1,\infty]$. A process $X\colon\mathbb{R}_{+}\to L^{p}(\mathcal{A},\mathbb{E})=L^{p}(\mathbb{E})$ is adapted if $X(t)\in L^{p}(\mathcal{A}_{t},\mathbb{E})$ for all $t\geq 0$. An $\boldsymbol{L^{p}}$-FV process is an adapted process $X\colon\mathbb{R}_{+}\to L^{p}(\mathbb{E})$ with locally bounded variation with respect to the noncommutative $L^{p}$ norm. By analogy with the classical notion of a semimartingale, we consider processes that can be decomposed as the sum of a martingale and an FV process: $X$ is a (continuous) $\boldsymbol{L^{p}}$-decomposable process if $X=X(0)+M+A$ for some $L^{p}$-continuous martingale $M\colon\mathbb{R}_{+}\to L^{p}(\mathbb{E})$ (Definition 3.11(ii)) and some $L^{p}$-continuous $L^{p}$-FV process $A\colon\mathbb{R}_{+}\to L^{p}(\mathbb{E})$ with $M(0)=A(0)=0$. In Section 3.2, we give many examples of $L^{p}$-decomposable processes and show that if $p\geq 2$, then the decomposition $X=X(0)+M+A$ is unique (Corollary 3.22). As in the classical case, we call $M$ the martingale part of $X$ and $A$ the FV part of $X$.
@@
We are now prepared to state our first main result: a noncommutative analog of Theorem 1.2.
@@
whenever $A,B\colon\mathbb{R}_{+}\to\mathcal{A}$ are adapted and continuous. These are special cases of integrals of trace biprocesses (Definition 3.7), which we introduce at the end of Section 3.1. The term “trace biprocess” is inspired by previous work on noncommutative stochastic calculus—specifically, that of Biane–Speicher [8]—in which integrals like $\int_{0}^{\boldsymbol{\cdot}}A(t)\,\mathrm{d}X(t)\,B(t)$ are treated by defining integrals of tensor-valued processes called “biprocesses.” Notably, however, integrals like $\int_{0}^{\boldsymbol{\cdot}}\mathbb{E}[A(t)\,\mathrm{d}X(t)]\,B(t)$ are not considered in previous work.
@@
The second main result is a construction of noncommutative quadratic covariation.
@@
Furthermore, $\lVert M(t)\rVert_{2}^{2}=\mathbb{E}\big[M^{*}(0)M(0)+\int_{0}^{t}\mathrm{d}M^{*}(s)\,\mathrm{d}M(s)\big]=\mathbb{E}\big[M(0)M^{*}(0)+\int_{0}^{t}\mathrm{d}M(s)\,\mathrm{d}M^{*}(s)\big]$.
@@
If $d,\ell\in\mathbb{N}$ and $f\colon\mathbb{R}^{d}\to\mathbb{C}^{\ell}$ is $k$-times continuously differentiable and $f_{*}(\mathbf{a})\coloneqq f\circ\mathbf{a}\in\mathcal{A}^{\ell}$ for $\mathbf{a}\in\mathcal{A}_{\operatorname{sa}}^{d}$, then $f_{*}\colon\mathcal{A}_{\operatorname{sa}}^{d}\to\mathcal{A}^{\ell}$ is an adapted $C^{k}$ map (Proposition 6.9 with $n=m=1$).
@@ 2.1 Noncommutative probability
$(L^{p}(\tau_{n}),\lVert\cdot\rVert_{L^{p}(\tau_{n})})=\Big(L^{p}(\Omega,\mathscr{F},P;\mathrm{M}_{n}(\mathbb{C})),\big\|\lVert\cdot\rVert_{L^{p}(\operatorname{tr}_{n})}\big\|_{L^{p}(\mathbb{E}_{P})}\Big)\qquad(1\leq p\leq\infty).$
@@ 2.2 Multilinear maps on $L^{p}$ spaces
We shall omit the subscripts when $k=1$. Finally, define $\mathbb{B}_{0}\coloneqq\mathcal{A}_{0+1}=\mathcal{A}_{1}$.
@@ 2.3 Trace $\ast$-polynomials
For example, $P(x_{1},x_{2},x_{3})=\operatorname{tr}(x_{1}x_{2}x_{1})x_{3}x_{2}^{2}x_{3}-\operatorname{tr}(x_{1})\operatorname{tr}(x_{2})x_{3}\in\mathrm{TrP}(x_{1},x_{2},x_{3})$, and sine $\operatorname{tr}$ is “tracial,” $P(x_{1},x_{2},x_{3})=\operatorname{tr}(x_{1}^{2}x_{2})x_{3}x_{2}^{2}x_{3}-\operatorname{tr}(x_{1})\operatorname{tr}(x_{2})x_{3}$ as well.
@@
  • $y_{j}^{*}$ appears precisely once in $P_{0}\cdots P_{\ell}$, and $y_{j}$ does not appear in $P_{0}\cdots P_{\ell}$.
@@ 3.1 Filtrations and adaptedness
For the rest of this section, write $(\mathcal{A},(\mathcal{A}_{t})_{t\geq 0},\mathbb{E})\coloneqq(\mathcal{A}_{1},(\mathcal{A}_{1,t})_{t\geq 0},\mathbb{E}_{1})$, and fix $k\in\mathbb{N}$. All tensor products below are over $\mathbb{C}$.
@@
    $_{k}(a_{1}\otimes\cdots\otimes a_{k+1})[b_{1},\ldots,b_{k}]=a_{1}b_{1}\cdots a_{k}b_{k}a_{k+1}\qquad(a_{i},b_{j}\in\mathcal{A}).$
@@
    $_{k}^{\mathbb{E}}(a_{1}\otimes\cdots\otimes a_{k+1})[b_{1},\ldots,b_{k}]=\mathbb{E}[a_{1}b_{1}\cdots a_{k}b_{k}]a_{k+1}\qquad(a_{i},b_{j}\in\mathcal{A}).$
@@ 4.1 Stieltjes integrals
Also, if $\Pi^{*}=(\Pi,\ast)\in\mathcal{P}_{I}^{*}$, then
$F^{\Pi^{*}}=F^{(\Pi,\ast)}\coloneqq 1_{\{a\}}F(a)+\sum_{s\in\Pi}1_{(s_{-},s]}F(s_{*})\colon I\to\mathcal{V}.$
If $t_{*}=\ast(t)$ is always the left endpoint $t_{-}$ of $t_{-},t]$, then $F^{\Pi}\coloneqq F^{(\Pi,\ast)}$.
The sets $\mathcal{P}_{I}$ and $\mathcal{P}_{I}^{*}$ are frequently directed by refinement. In this paper, we shall instead direct them by mesh $|\cdot|$, i.e., $\Pi\leq\Pi^{\prime}$ and $(\Pi,\ast)\leq(\Pi^{\prime},\ast^{\prime})$ whenever $|\Pi|\geq|\Pi^{\prime}|$. Here is an elementary clarification of what convergence of nets directed by partition mesh means. We leave the proof to the reader.
@@
    If $H$ is continuous and $\mathcal{V}$ is normed, then $H^{\Pi^{*}}\to H$ uniformly on compact subsets of $I$ as $|\Pi|\to 0$.
As the reader may easily verify, it suffices to treat the $b<\infty$ case. Write $\iota\colon I\to\mathbb{R}$ for the inclusion. Observe that if $\Pi^{*}=(\Pi,\ast)\in\mathcal{P}_{I}^{*}$ and $t\in(a,b]$, then $H^{\Pi^{*}}=H\circ\iota^{\Pi^{*}}$, $0<t-\iota^{\Pi}(t)\leq|\Pi|$, and $\big|t-\iota^{\Pi^{*}}(t)\big|\leq|\Pi|$. Using these observations, the desired results follow easily from the definitions and, in the case of (ii), the fact that continuous functions $I\to\mathcal{V}$ are uniformly continuous. ∎
@@
$\sum_{t\in\Pi}H(t_{*})[F(t\wedge\cdot)-F(t_{-}\wedge\cdot)]\xrightarrow[\Pi^{*}\in\mathcal{P}_{I}^{*}]{|\Pi|\to 0}\int_{a}^{\boldsymbol{\cdot}}H(t)[\mathrm{d}F(t)]$
@@
In particular, the integral map $\mathbb{S}_{I}\ni S\mapsto\int_{a}^{\cdot}S[\mathrm{d}\mu_{F}^{0}]\in\ell^{\infty}(I;\mathcal{W})$ extends uniquely to a bounded linear map $J\colon\overline{\mathbb{S}_{I}}\to\ell^{\infty}(I;\mathcal{W})$, where $\overline{\mathbb{S}_{I}}$ is the closure of $\mathbb{S}_{I}$ in $\ell^{\infty}(I;B(\mathcal{V};\mathcal{W}))$. Now, if $H\colon I\to B(\mathcal{V};\mathcal{W})$ is continuous and $\Pi^{*}\in\mathcal{P}_{I}^{*}$, then $H^{\Pi^{*}}\in\mathbb{S}_{I}$, and by Lemma 4.4(ii), $H^{\Pi^{*}}\to H$ in $\ell^{\infty}(I;B(\mathcal{V};\mathcal{W}))$ as $|\Pi|\to 0$. Thus, $H\in\overline{\mathbb{S}_{I}}$, and by the continuity of $J$,
$\sum_{t\in\Pi}H(t_{*})[F(t\wedge\cdot)-F(t_{-}\wedge\cdot)]=\int_{a}^{\cdot}H^{\Pi^{*}}\big[\mathrm{d}\mu_{F}^{0}\big]=J\big[H^{\Pi^{*}}\big]\xrightarrow[\Pi^{*}\in\mathcal{P}_{I}^{*}]{|\Pi|\to 0}J[H]$